OpenStudy (moonlight93):

I have a quiz tomorrow , please help me out with this one, similar problem might be on the quiz. A 200 KG block slides down a frictionless slope. the slope is 30m long and the block slides from an elevation of 10m. Using the principle of conversion of energy, how fast is the block going when it hits the bottom? Note: Ignore the dimensions of the mass

OpenStudy (anonymous):

ok so first u start with knowing that it has only potential gravitational energy at the top of the slope, and only kinetic at the bottom, so G=K, which is just mgh=1/2mv^2.The length of the slope doesnt really matter. Technically the mass cancels out on both sides but just to be safe you can leave it in: 200(9.8)(10)=1/2(200)v^2, which is 19600=100v^2, and once u divide both sides by 100, u can square root both sides and get v=14 :^)

OpenStudy (moonlight93):

oh, that doesn't look so bad after all, thank you sooooooo much for helping :)